40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16549     Entry time: Thu Jan 6 15:10:38 2022
Author: Koji 
Type: Update 
Category: SUS 
Subject: ITMX Chamber work 

[Anchal, Koji]

=== Summary ===
- ITMX SD OSEM migration done
- LO1 OSEM insertion and precise adjustment (part 1) done
- LO1 POS/PIT/YAW/SD motions were damped


=== General Remarks ===
- 15:00 Entered into ITMX.
- We were equipped with N95 and took physical distance as much as possible.
- 17:00 Temporarily came out from the lab.
- 18:30? Came into the chamber again
- 20:00 Sus damped. OSEM work continues
- 21:00 OSEM installation work done. Exit.

=== ITMX SD OSEM position swap ===
- Moved the LO1 suspension to the center of the chamber
- Removed the ITMX SD OSEM from the right side (west side) and tried to move it to the other side.
- Noted that the open light output of the ITMX SD was 908 at the output of the SDSEN filter module. So the half-light target is 454. These numbers include the "cnt2um" calibration of 0.36. That means the open light raw ADC count was supposed to be 2522.

- The OSEM set screw (silver plated, with a plunger) was removed from the old position. We first tried to recycle it to the other side, but it didn't go into the thread with fingers. After making ourselves convinced that the threaded hole was identical for both sides, we decided to put the new identical plunger set screw with an Allen-key was used to put it in and it went in!
- Now the ITMX SD OSEM was inserted from the east side. Once we saw some shadow on the OSEM signal, the SD damping was turned on with the previous setting. And this successfully damped the side motion. ⭕️
- A bit finer adjustment has been done. After a few trials, we reached the stable output of ~400. Considering the temporary leveling of the table, we decided this is enough for now ⭕️. The set screw was tightened.
- To make the further work safer w.r.t the ITMX magnets, Anchal fastened the EQ stops of the ITMX sus except for the bottom four.
- Photo: [Attachment 1]

=== LO1 OSEM installation ~ wiring ===
- Now LO1 was moved back to the planned position.
- For the wiring, we (temporarily) clamped the in-vac DSUB cables to the stack table with metal clamps.
- Started plugging the OSEMs into the DSUB cables.
- Looking at the LO1-1 cable from the mating side with the longer side top: The top-right pin of the female connector is Pin1 as usual. From right to left LL / UR / UL coils were inserted one by one while looking at the OSEM PD signals.
- LO1-2 cable has the LR / SD coils (from the right to the left) were connected.
- Photo: [Attachment 2]

- LO1 Open light levels (raw ADC counts) the 2nd number is the target half-light level

  • UL 27679 (-> 13840)
  • UR 29395 (-> 14697)
  • LR 30514 (-> 15257)
  • LL 27996 (-> 13998)
  • SD 26034 (-> 13017)

=== RTS Filter implementation ===

- Anchal copied the filter module settings from other suspensions.
- We also implemented the simple input and output matrices.

=== LO1 OSEM insertion ===

- We struggled to make the suspension freely swinging with the OSEMs inserted.

- It seemed that the magnets were sucked to the OSEMs due to magnetic components.
- It turned out that the OSEMs were not fastened well and not seated in the holder plates.
- Once this was fixeded, we found that the mirror height is too high for the given OSEM heights.
  The suspension height (or the OSEM height should be decided with the OSEMs not inserted but fully fastened to prevent misalignment of them.

- Decided to lift up the OSEM plates in situ.
- Soon we found that the OSEM holder plates are not fastened at all [Attachment 3 arrows]
- The plates were successfully lifted up and
the suspension became much more freely swinging even with the OSEMs inserted. ⭕️

=== LO1 damping and more precise OSEM insertion ===

- Once the OSEMs were inserted to the light level of 30~70%, we started to try to dampen the motion. The side damping was somewhat successful, but the face ones were not.
- We checked the filters and found the coil output filters didn't have the alternating signs.
- Once the coil signs were corrected, the damping became more straight forward.
- And the robust damping allowed us the fine-tuning of the OSEM insertion.

- In the end, what we had for the light levels were

  • UL 14379 (52%)
  • UR 14214 (48%)
  • LR 14212 (47%)
  • LL 12869 (46%)
  • SD 14358 (55%)

The damping is working well. [Attachment 4]


Post continues at 40m/16552.

Attachment 1: PXL_20220107_044739280.MP.jpg  6.768 MB  Uploaded Thu Jan 6 21:33:36 2022  | Show | Hide all | Show all
Attachment 2: PXL_20220107_044958224.jpg  1.674 MB  Uploaded Thu Jan 6 21:37:03 2022  | Hide | Hide all | Show all
PXL_20220107_044958224.jpg
Attachment 3: PXL_20220107_044805503.NIGHT.jpg  1.347 MB  Uploaded Thu Jan 6 21:38:38 2022  | Hide | Hide all | Show all
PXL_20220107_044805503.NIGHT.jpg
Attachment 4: Screen_Shot_2022-01-06_at_20.54.04.png  243 kB  Uploaded Thu Jan 6 22:31:23 2022  | Hide | Hide all | Show all
Screen_Shot_2022-01-06_at_20.54.04.png
ELOG V3.1.3-